# Enhancing Layer Attention Efficiency through Pruning Redundant Retrievals

### v3

@@ Abstract
Growing evidence suggests that layer attention mechanisms, which enhance interaction among layers in deep neural networks, have significantly advanced network architectures. However, existing layer attention methods suffer from redundancy, as attention weights learned by adjacent layers often become highly similar. This redundancy causes multiple layers to extract nearly identical features, reducing the model’s representational capacity and increasing training time. To address this issue, we propose a novel approach to quantify redundancy by leveraging the Kullback-Leibler (KL) divergence between adjacent layers. Additionally, we introduce an Enhanced Beta Quantile Mapping (EBQM) method that accurately identifies and skips redundant layers, thereby maintaining model stability. Our proposed Efficient Layer Attention (ELA) architecture, improves both training efficiency and overall performance, achieving a 30% reduction in training time while enhancing performance in tasks such as image classification and object detection.
@@ 1 Introduction
Numerous studies have successfully demonstrated that enhancing the representational ability of deep convolutional neural networks (DCNNs) significantly improves their performance across various tasks. For instance, ResNet He et al. (2016) introduced skip connections to facilitate gradient flow, effectively mitigating the problem of performance degradation in deep networks. DenseNet Huang et al. (2017) further enhanced model capacity by reusing information from all previous layers. In addition to these innovations, attention mechanisms have further enhanced model performance. Channel attention mechanisms Hu et al. (2018); Wang et al. (2020b); Qin et al. (2021); Ouyang et al. (2023), refine feature selection by prioritizing significant channels, allowing networks to focus on the most relevant feature maps. Complementing this, spatial attention mechanisms Woo et al. (2018); Wang et al. (2018); Li et al. (2023a); Wang et al. (2024c), optimize performance by directing attention to critical spatial regions, improving the network’s ability to capture spatial hierarchies and refine learned representations.
@@
Recently, enhancing interactions between hierarchical layers has emerged as another effective approach to improving model’s representational ability. DIANet Huang et al. (2020) incorporated a modified LSTM block along the depth of the network to promote the interactions between layers. Similarly, RLANet Zhao et al. (2021) employs a lightweight recurrent aggregation module to effectively reuse information from previous layers. BViT Li et al. (2023b) processes the output of each Transformer layer in the ViT Dosovitskiy (2020) architecture through attention mechanism, effectively enhancing model’s representational ability and overall performance. To further enhance cross-layer interaction, MRLA Fang et al. (2023) first introduced layer attention, leveraging a multi-head attention mechanisms across layers to align features from current layer with those of all preceding layers.
However, we have identified a significant drawback in existing layer attention mechanisms: the inherent redundancy within the attention process across layers. As shown in Figure 2(a), the attention weights learned by three consecutive layers (the 4th, 5th, and 6th layers) are nearly identical, with the attention focusing on just one layer. This suggests that these layers are extracting almost identical features, which undermines the model’s representational capacity. For instance, in the MRLA-B model, as the network depth increases, accuracy drops noticeably, with performance even falling below that of the ResNet baseline, as depicted in Figure 2(b). The lightweight layer attention network MRLA-L achieves impressive performance, but continues to struggle with low training efficiency induced by redundancy. Therefore, eliminating redundant attention layers from the layer attention network is essential to maintain both model performance and training efficiency, as demonstrated in Figure 1.
To assess redundancy in layer attention networks, we propose using KL divergence to evaluate the similarity of attention weights between adjacent layers. In a layer attention network, the attention weights of each layer can be treated as a probability distribution, and KL divergence measures the difference between these distributions. Specifically, a smaller KL divergence value indicates greater functional similarity between adjacent layers, as shown in Figure 3. Our findings reveal that the distribution of these KL divergences remains stable throughout training. However, some KL divergences may become overly concentrated, making it difficult to accurately identify which layers to prune. To address this, we introduce an innovative algorithm, Enhanced Beta Quantile Mapping (EBQM), which make the KL divergence of distributions suitable for pruning layers. This improves the stability of the model’s performance. Experimental results validate the effectiveness of our method, showing improvements in both speed and accuracy over state-of-the-art techniques across various tasks.
@@
We propose a novel approach that employs KL divergence to assess redundancy in layer attention networks, which quantifies the similarity of attention weights between adjacent layers.
An innovative algorithm, named EBQM, is introduced to skip redundant attention layers while maintaining model stability.
Our ELA architecture successfully reduces redundancy in existing layer attention mechanisms, resulting in an approximate 30% increase in training speed and improved performance on various tasks, including image recognition and object detection.
@@ 2.1 Layer Interaction
A growing number of research has focused on layer interaction in deep convolutional neural networks. DIANet Huang et al. (2020) incorporated a parameter-shared DIA block across all layers to facilitate information flow between layers. Similarly, RLANet Zhao et al. (2021) proposed a lightweight recurrent aggregation module that efficiently reuses information from previous layers to improve feature extraction in subsequent layers. Zhao et al. (2022) integrated information from the preceding layers with the current layer through a dot product operation, significantly improving the model performance. BViT Li et al. (2023b) applied an attention mechanism to the output of each layer in Vision Transformer (ViT), effectively integrating information throughout the model. ST-Unet Zhang et al. (2023) proposed a feature enhancement module during the concatenation of the inter-layer feature map in U-Net Ronneberger et al. (2015), which improves the feature extraction process in each layer. MRLA Fang et al. (2023) treated the features of each layer as tokens, employing an attention mechanism to promote interaction between hierarchical layers, further reinforcing the connectivity between layers. DLANet Wang et al. (2024b) employed a modified LSTM module at every layer, ensuring dynamic information update before MRLA attention operation.
@@ 2.2 Efficient Attention
Traditional attention mechanisms are often hindered by the complexity, which limits their scalability. Multiple studies, such as Shen et al. (2021); Ma et al. (2021), addressed this issue by utilizing linear attention structures to reduce the computational complexity. Furthermore, Li et al. (2022) accelerated attention operations in both NLP and vision tasks by pruning tokens with small attention scores. Guo et al. (2023) mitigated the issue of Transformers overemphasizing specific tokens by restricting the attention mechanism to the local neighborhood of each token. SKIPAT Venkataramanan et al. (2023) identified the phenomenon of functional similarity in successive layers within vision transformers (ViT) and proposed replacing these redundant transformer layers with a parametric function to enhance efficiency. Similarly, EfficientViT Liu et al. (2023) employed cascaded group attention to improve attention diversity in ViT, effectively addressing redundancy.Additionally, Wang et al. (2024a) improved efficiency in ViT by pruning less important or similar tokens from the attention map.
@@ 3 Method
We will begin by recalling the existing layer attention architecture and highlighting its redundancy. Next, we will explain how KL divergence can be used to identify redundant attention layers. Finally, we will explain the issue of overly concentrated KL divergences and present our Enhanced Beta Quantile Mapping Algorithm, which resolves this problem and stabilizes model performance.
@@ 3.1 Revisiting Layer Attention
Layer attention Fang et al. (2023) is a novel mechanism that enhances inter-layer dependencies by allowing each layer in a neural network to retrieve query-relevant information from all preceding layers. Given as the feature of layer , the layer attention mechanism is employed as the following formulas:
The function serves to extract Query from the current layer , while the functions and are responsible for extracting Key and Value matrix from , where denotes the -th layer feature. Then the layer attention can be calculated as:
This represents the structure of MRLA-B defined by Fang et al. (2023). For clarity, the softmax and the scaling factor in the multihead attention mechanism defined by Vaswani (2017) are omitted here. By substituting :
@@
In this case, is defined as a learnable vector to dynamically adjust the contribution of prior layers during the attention process.
@@ 3.2 Motivation
MRLA leverages a multihead attention mechanism to strengthen inter-layer interactions. However, a key limitation of MRLA is the redundancy observed within the network, where the attention weights learned by certain adjacent attention layers are very similar. This similarity indicates that these layers extract nearly identical features. To further investigate this phenomenon, we utilized the Score-CAM Wang et al. (2020a) method to visualize the outputs of sequential MRLA layers, as shown in Figure 3. The figure presents the feature map visualizations of successive MRLA layers in ResNet-101, using images from the ImageNet-1K dataset. The results demonstrate that when the KL divergence between the attention weights of two MRLA layers is small, their feature maps are remarkably similar, highlighting their functional redundancy. Moreover, the number of such redundant layers increases with network depth, leading to a gradual decline in the model’s representation capability and an unnecessary increase in training time.
The presence of redundant attention layers undermines the capability of the original backbone network, as numerous layers end up performing identical tasks. Skipping these redundant attention layers while retaining the original backbone layers allows the network to both leverage the layer attention mechanism to retrieve information from previous layers and ensure feature diversity across layers. This approach enhances the model’s capacity and reduces training time without introducing additional parameters. Therefore, this paper aims to develop an innovative method to identify and precisely skip these redundant layers.
3.3 Layer Pruning Based on KL Divergence
@@
KL Divergence In our approach, we employ the KL divergence to measure the difference between probability distributions derived from the attention weights of adjacent layers in a neural network. The KL divergence between two discrete probability distributions and is defined as:
Here, and represent the probabilities from distributions and , respectively. The KL divergence quantifies the difference between these two distributions, where a smaller KL divergence indicates that the two distributions are more similar, while a larger value reflects greater dissimilarity.
KL Divergence in Layer Attention Let the output of the -th layer be denoted as , where , , and represent the number of channels, height, and width, respectively. The output can be expressed as a linear combination of features:
where is the -th value in , and represents the distribution of attention weights.
Similarly, the output of the -th layer, , is given by:
@@
where is another probability distribution derived from the attention mechanism at the -th layer. Notably, the distribution has one additional value compared to .
To ensure the two distributions are comparable, we append an infinitesimally small value (where ) to the distribution so that it matches the dimensionality of , i.e.,
We then compute the KL divergence between these distributions using the standard formula:
@@
In a layer attention network, the feature maps remain consistent between layers and , which means that the difference between the outputs of the two layers is entirely determined by the difference in their attention weight distributions and . The KL divergence between and thus directly reflects the similarity between the outputs and , as demonstrated in Figure 3.
@@
Attention-Layer Pruning Algorithm A training process is divided into several stages, and in each stage, we begin by calculating the KL divergences between the attention weights of adjacent MRLA layers. Here we do not explicitly define the number of layers to skip. This is because, at each training stage (except the initial one), the KL divergence distribution is influenced by the results of the previous stage, which affects the number of layers to be discarded. If the same number of layers is removed from each stage, some critical attention layers in certain stages may be discarded, which can degrade the model’s performance, as demonstrated in the results of Table 4. However, within each individual training stage, the distribution of KL divergence remains relatively stable.
@@
The proposed algorithm for layer pruning, as detailed in Algorithm 1, uses a threshold to determine which layers to discard. Initially, we focus on KL divergence values below the computed -quantile and normalize these values to the range for ease of threshold configuration. However, in certain cases, the normalized KL divergence values tend to cluster within a narrow range. If a single threshold is applied, even small fluctuations in KL divergence values across different training runs or stages can lead to significant variations in the number of skipped MRLA layers. For instance, if the threshold is set to 0.3, six layers with KL divergence values ranging from 0.31 to 0.33 in one training run may fall within the range of 0.27 to 0.29 in another. This discrepancy results in a six-layer difference in the layers skipped between the two runs. Although limiting the normalization process to KL divergence values below the -quantile alleviates the issue to some extent, it still has a substantial impact on model performance and robustness, as shown in Table 4.
@@
To address this issue, we propose preprocessing the distribution of normalized KL divergence values prior to thresholding. Inspired by the Quantile Mapping Algorithm, a technique widely applied in meteorology and hydrology to align disorganized data with a target distribution, we propose an innovative Enhanced Beta Quantile Mapping (EBQM) Algorithm to reshape the distribution of values. Here, we reshape the data using the cumulative distribution function (CDF) of the Beta distribution.
The CDF of a Beta distribution can be illustrated in mathematical terms:
on the interval , is given by:
where and are shape parameters, and is the Beta function, expressed as:
Unlike other distribution functions, such as the Gamma distribution, the CDF of the Beta distribution is uniquely defined within the interval for both its input and output. This property aligns perfectly with the normalized KL divergence values we aim to reshape, as they also fall within the range. Consequently, the normalized KL divergence values can be directly utilized as inputs to the Beta distribution’s CDF for efficient and precise distribution reshaping. By applying this Beta-distribution-based transformation before thresholding, and adjusting the parameters and , we can effectively disperse clustered values under almost all conditions.

### v4

@@ Abstract
Growing evidence highlights the impact of layer attention mechanisms in deep neural networks, which enhance inter-layer interactions and representational ability. However, existing layer attention methods suffer from redundancy, as adjacent layers often learn highly similar attention weights. This redundancy leads to multiple layers retrieving nearly identical features, diminishing the model’s representational capacity and increasing training time. To address this issue, we propose an Efficient Layer Attention (ELA) architecture composing two key components: first, we quantify redundancy by leveraging Kullback-Leibler (KL) divergence between adjacent layers; second, we introduce an Enhanced Beta Quantile Mapping (EBQM) algorithm to accurately detect and skip redundant retrieving, ensuring model stability while boosting efficiency. Experimental results demonstrate that ELA not only enhances overall performance but also reduces training time by 30%, yielding improvements in tasks such as image classification and object detection.
@@ 1 Introduction
Numerous studies have demonstrated that enhancing inter-layer interactions in deep convolutional neural networks (DCNNs) can substantially improve performance across various tasks. For instance, ResNet [10] introduced skip connections to facilitate gradient flow, effectively mitigating the issue of performance degradation in very deep networks. Building upon this idea, DenseNet [15] further strengthened layer interactions by reusing information from all previous layers. In addition to these innovations, attention mechanisms have also played a crucial role in boosting model performance. Channel attention mechanisms [13, 34, 27, 23], refine feature selection by prioritizing the most informative channels, enabling the network to focus on the salient feature maps. Complementarily, spatial attention mechanisms [37, 35, 17, 36] improve the network’s ability to model spatial hierarchies by directing attention to the most relevant spatial regions, thereby refining the learned representations.
@@
Recently, enhancing inter-layer interactions through layer attention mechanisms has emerged as an effective strategy for improving model’s representational ability. DIANet [16] introduced an improved LSTM block along the depth of the network to facilitate interactions between layers. Similarly, RLANet [41] employs a lightweight recurrent aggregation module to effectively reuse information from previous layers. BViT [18] further aggregates the outputs of each Transformer layer in the ViT [7] architecture through attention mechanism, significantly boosting overall performance. MUDD [38] enhances layer interaction through dynamic connection weights, which depends on each sequence position in a Transformer block. To further enhance cross-layer interaction, MRLA [8] introduced layer attention that leverages a multi-head attention across layers to align features from current layer with those from all preceding layers. Building upon this, Wang et al. argued that MRLA operates in a static manner, which limits its efficiency. To address this, they proposed Dynamic Layer Attention (DLA) [33], which restores the dynamic context representation for more efficient layer attention.
However, we have identified a significant drawback in existing layer attention mechanisms: the redundant retrieval for previous feature across layers. As shown in Figure 2(a), the attention weights learned by three consecutive layers (the 4th, 5th, and 6th) are nearly identical, all heavily focusing on the 2nd layer. This consistent retrieval pattern is not merely a coincidence—it indicates a systemic redundancy in the attention mechanism. When multiple layers repeatedly attend to the same previous features, the model wastes computation and fails to fully exploit the expressive potential of deeper layers. This issue becomes particularly evident in practice. As shown in Figure 2(b), when the network depth exceeds 56 layers, the performance of MRLA-B declines as the depth increases, and with ResNet-110 as the backbone, its accuracy even drops below that of the original ResNet-110 baseline. Although MRLA-L maintains performance gains as the network deepens, the improvement slows down, primarily due to the reduced training efficiency caused by redundant attention patterns.
Based on these observations, we propose a key research question: Is it necessary to relearn similar attention patterns in each layer? Instead of repeatedly retrieving the same earlier features, later layers—such as the 5th and 6th—can skip the layer attention operation entirely when redundancy is detected, since the relevant information has already been retrieved by the preceding layer. This not only reduces computational overhead, but also alleviates the issue of over-representing certain features, leading to more efficient and balanced information flow across the network.
To address this issue, we propose an Efficient Layer Attention (ELA) architecture consisting of two key components: ❶ a KL divergence-based strategy to quantify attention redundancy, and ❷ an Enhanced Beta Quantile Mapping (EBQM) algorithm to guide the pruning of redundant retrievals. Specifically, we treat attention weights in each layer as probability distributions and compute the KL divergence between successive layers. A low KL divergence indicates that the attention behavior of the current layer is highly similar to that of the previous layer, suggesting that we can directly discard the retrieval from current layer. However, KL divergence may occasionally fluctuate during training, making it unreliable for pruning in isolation. To address this, EBQM calibrates these divergences into a stable mapping that enables reliable identification of redundant layers. Together, these components allow ELA to reduce unnecessary computation while maintaining or even improving model accuracy.
@@
We introduce a novel method that uses KL divergence to quantify redundancy in layer attention networks, enabling the identification of similar attention patterns across consecutive layers.
We propose an innovative algorithm, Enhanced Beta Quantile Mapping (EBQM), which effectively prunes redundant attention retrievals while preserving model stability and performance.
We present the Efficient Layer Attention (ELA) architecture, which significantly reduces computational overhead by reusing attention weights across layers, resulting in approximately 30% faster training and improved performance across tasks such as image recognition and object detection.
@@ 2.1 Layer Interaction
A growing number of research has focused on enhancing layer interaction in deep convolutional neural networks. DIANet [16] introduced a parameter-shared DIA block applied across all layers to facilitate information flow throughout the network. Similarly, RLANet [41] proposed a lightweight recurrent aggregation module that efficiently reuses information from previous layers to enhance feature extraction in subsequent ones. [42] integrated features from preceding layers with the current layer using a dot product operation, significantly boosting the model performance. BViT [18] applied an attention mechanism to the output of each layer in Vision Transformer (ViT), enabling effective integration of information across layers. ST-Unet [40] introduced a feature enhancement module at the inter-layer concatenation stage in U-Net [28], thereby improving feature extraction at each level. MRLA [8] treated features from each layer as tokens and employed an attention mechanism to promote interaction between hierarchical layers, further reinforcing the inter-layer connectivity. DLANet [33] employed a modified LSTM module at each layer to dynamically update information prior to the MRLA attention operation. Finally, MUDD [38] enhanced inter-layer interaction by dynamically adjusting connection weights based on each token’s position within a Transformer block.
@@ 2.2 Efficient Attention
Traditional attention mechanisms are often constrained by their quadratic complexity , which limits their scalability. To address this limitation, several studies have proposed more efficient alternatives. For example, [29, 22] introduced linear attention structures that significantly reduce computational overhead. Furthermore, [19] accelerated attention operations in both NLP and vision tasks by pruning tokens with low attention scores. [9] mitigated the issue of Transformers overemphasizing specific tokens by restricting the attention mechanism to a local neighborhood. SKIPAT [30] observed functional similarity across successive layers in Vision Transformers (ViTs) and proposed replacing redundant transformer blocks with lightweight parametric functions to improve efficiency. Similarly, EfficientViT [21] employed cascaded group attention to enhance attention diversity and reduce redundancy in ViTs. Additionally, [32] improved computational efficiency by pruning less important or similar tokens from the attention map.
@@ 3 Method
We start by revisiting the layer attention mechanism, followed by the motivation for our proposed Efficient Layer Attention (ELA) architecture. KL divergence is roughly employed to measure the redundant retrievals when adjacent layers attend to similar earlier layers. Then, we propose the Enhanced Beta Quantile Mapping (EBQM) algorithm, which effectively smooths the KL divergence distribution to detect and skip redundant retrievals.
@@ 3.1 Revisiting Layer Attention
Layer attention [8] enhances inter-layer interactions by employing attention mechanisms, enabling each layer to selectively retrieve relevant information from all preceding layers. Given that represents the feature of -th layer, where , , and denote the number of channels, height, and width, respectively, we first extract the query, key, and value representations as follows:
Here, extracts the query from , while and extract the key and the value from the -th layer, respectively. The full key and value matrices and are constructed by concatenating the representations from all preceding layers up to . Then the layer attention can be calculated as:
Where represents the -th attention output. For clarity, the softmax and the scaling factor are omitted here. By substituting:
@@
Where is a learnable vector to dynamically adjust the contribution of prior attention output.
@@ 3.2 Motivation
In our preliminary experiments (Figure 2(a)), we observed that adjacent layers in layer attention networks often learn highly similar attention weights. Since layer attention retrieves information from previous layers based on these weights, such similarity leads to nearly identical attention outputs across consecutive layers. To further validate this, we employed Score-CAM [31] to visualize the attention outputs in MRLA-101. As illustrated in Figure 3, when the KL divergence between adjacent layers’ attention distributions is low, their attention outputs also exhibit high visual similarity. This indicates redundant retrieval behavior—repeatedly attending to the same source of information—which not only introduces unnecessary computation but also reduces the diversity of learned features, limiting the model’s representational capacity.
Motivated by these observations, we propose a method that progressively learns to skip redundant retrievals during training. Instead of recomputing attention in every layer and discarding it afterward based on KL divergence, our approach integrates redundancy detection into training itself. When the KL divergence between the current layer and the previous layer remains consistently small, the model learns to skip the attention retrieval step in subsequent layers, effectively pruning these operations from the architecture. This reduces computational overhead while improving training efficiency and maintaining expressive power.
3.3 Preliminary Knowledge
@@
KL Divergence We first utilize the Kullback–Leibler (KL) divergence to roughly approximate the similarity between attention weights learned in adjacent layers of a neural network. Given two discrete probability distributions and , the KL divergence is defined as:
The KL divergence measures the difference between two distributions: a smaller value implies higher similarity, while a larger value indicates greater divergence.
KL Divergence in Layer Attention The output of the -th layer can be expressed as a linear combination of value vectors:
@@
where is the -th value in , and denotes the attention weight from -th layer, which can be interpreted as a probability distribution.
Similarly, the output of the -th layer can be written as:
where is the attention weight derived from -th layer. Notably, the distribution contains one additional term, , compared to .
To enable a valid comparison between the two distributions, we pad with an infinitesimal value (where ), resulting in , so that its dimensionality matches that of .
The KL divergence between the two distributions is then computed as:
@@
In a layer attention network, the previous values remain consistent between layers and . Consequently, the difference between the outputs and is solely attributed to the change in their corresponding attention weight distributions, and . Therefore, the KL divergence between and directly captures the degree of similarity between the two outputs.
3.4 Efficient Layer Attention
ELA Architecture We present the details of our proposed Efficient Layer Attention (ELA) architecture (Shown in Figure 1). Firstly, we compute the KL divergences between attention weights learned from all adjacent layers using Equation (11), and collect the results into a divergence vector . This vector is then passed through the EBQM module , which smooths the divergence values as follows:
Next, we apply a pruning mechanism based on a pre-defined threshold . A binary mask is generated as:
This mask determines whether the attention outputs from each layer should be preserved or pruned. In each layer , the representations of query, key, and value are then updated as follows:
Then refined query and are then used to compute the layer attention for layer as described in Equation (5).
@@ 3.4 Efficient Layer Attention
Training Strategies Our training process is divided into multiple stages. At the beginning of each stage, we compute the KL divergences between the attention weights of adjacent MRLA layers. Notably, we do not predefine a fixed number of layers to skip. This is because the KL divergence distribution at each stage (except the initial one) is influenced by the pruning results from previous stages, which dynamically affects the number of layers to be discarded. If a fixed number of layers were removed at each stage, important attention layers might be pruned in some stages, ultimately degrading model performance—as evidenced by the results in Table 4. However, within each stage, the KL divergence distribution remains relatively stable.
3.5 EBQM Algorithm
@@
The proposed pruning approach, detailed in Algorithm 1, leverages a threshold to determine which layers to discard. Initially, we focus on the KL divergence values that fall below the -quantile and normalize these values to the range to simplify thresholding. However, these normalized values are often tightly clustered, making the pruning highly sensitive to small variations across training runs. For example, with a threshold , six layers with KL divergences between 0.31 and 0.33 in one run may shift to between 0.27 and 0.29 in another. This results in up to six fewer skipped layers in the latter case, introducing inconsistencies that can significantly impact model performance and robustness, as shown in Table 4.
@@
To mitigate this issue, we propose preprocessing the normalized KL divergence distribution before thresholding. Drawing inspiration from the Quantile Mapping Algorithm—a technique widely used in meteorology and hydrology to adjust data distributions—we introduce the Enhanced Beta Quantile Mapping (EBQM) algorithm. EBQM reshapes the distribution using the cumulative distribution function (CDF) of the Beta distribution.
The CDF of the Beta distribution is defined as:
where is given by:
and is the Beta function:
Unlike other distributions such as the Gamma distribution, the Beta distribution’s CDF is strictly defined within the interval for both its input and output. This makes it especially suitable for transforming normalized KL divergence values, which also lie within . By applying this transformation and tuning the shape parameters and , we can effectively spread out tightly clustered values, enabling more stable and robust pruning decisions across different training stages.